OpenStudy (anonymous):

a circle has the equation x^2+y^2=16. find the center, radius and graph

OpenStudy (anonymous):

how do I put this in standard form and general form pls?

OpenStudy (anonymous):

\[\LARGE (x-0)^2+(y-0)^2=4^2\] \[\LARGE Q(0,0) \quad r=4\]

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 (a,b) = Centre r = Radius Centre is (0,0) and Radius is \[\sqrt{16} = 4\]

OpenStudy (anonymous):

thank you ! both of you!

OpenStudy (anonymous):

Although this looks like ellipse, is a circle with radius 4 :)

OpenStudy (anonymous):

and it is an ellipse, as well as a circle.